THE JAHN-TELLER EFFECT IN NO_{3}: SHEDDING NEW LIGHT ON THE DARK Ã ^{2}E′′ STATE

ANDREI DEEV, KANA TAKEMATSU, DAVID J. ROBICHAUD, \underline{MITCHIO OKUMURA}, Arthur Amos Noyes Laboratory of Chemical Physics, California Institute of Technology, Pasadena CA 91125; JOHN F. STANTON, Department of Chemistry, University of Texas, Austin, TX, 78712.

The Jahn-Teller Effect in the lowest excited states of the nitrate radical NO_{3} remains poorly understood. In this paper, we examine the first excited state, Ã ^{2}E′′, in a joint experimental and theoretical study. The forbidden Ã ^{2}E′′ ← X̃ ^{2}A′_{2} transition is recorded by Near Infrared Cavity Ringdown Spectroscopy (CRDS) at medium resolution from 6000 to >10,000 cm^{−1}, extending and refining our preliminary results. We observe over 30 major vibronic bands, most with resolvable rotational structure. Tentative assignments are made on the basis of high level EOMIP-CCSD(T) calculations of the vibronic Hamiltonian, which includes up to quartic terms. Our results indicate that the Ã state undergoes static Jahn-Teller distortion, with strong vibronic coupling among both degenerate modes and the symmetric stretch. The Jahn-Teller effect in NO_{3} appears to be complex and unusual, and remains a difficult but fascinating challenge.